If 2 less than three-fourths of a number is 17 less than one-eighth of the number, find the number
Answer by Cromlix(4381)

Hi there,
Make the unknown number = x
3/4 x - 2 = 1/8 x - 17
Collect like terms
3/4x - 1/8x = -17 + 2
Lowest denominator for x is 1/8
(6 - 1)/8 x = -15
5/8x = -15
Multiply both sides by 8/5
x = -24
Number is -24.
3/4(-24) - 2 = 1/8(-24) -17
-18 - 2 = -3 - 17
-20 = -20
Hope this helps :-)